### 18 Dec 2023 v3

# An alternative form of supersymmetry with reduced cross-sections and modified experimental signatures

Abstract
There is a convincing case for some form of supersymmetry, but conventional supersymmetry (SUSY) has been plagued by many unsolved theoretical difficulties since its inception half a century ago. Even more importantly, not a single SUSY superpartner has been observed up to surprisingly high experimental limits. These failures suggest that it is appropriate to rethink the meaning of supersymmetry at the most fundamental level. Here we consider a radically different form of supersymmetry (called susy here to avoid confusion), which initially combines standard Weyl fermion fields and primitive (unphysical) boson fields. A stable vacuum then requires that the initial boson fields, whose excitations would have negative energy, be transformed into three kinds of scalar-boson fields: the usual complex fields $\phi$, auxiliary fields $F$, and real fields $\varphi$ of a new kind (with degrees of freedom and gauge invariance preserved under the transformation). The requirement of a stable vacuum thus imposes Lorentz invariance, and also immediately breaks the initial susy – whereas the breaking of conventional SUSY has long been a formidable difficulty. Even more importantly, for future experimental success, the present formulation may explain why no superpartners have yet been identified: Embedded in an $SO(10)$ grand-unified description, most of the conventional processes for production, decay, and detection of sfermions are excluded, and the same is true for many processes involving gauginos and higgsinos. This implies that superpartners with masses $\sim 1$ TeV may exist, but with reduced cross-sections and modified experimental signatures. For example, a top squark (as redefined here) will not decay at all, but can radiate pairs of gauge bosons and will also leave straight tracks through second-order (electromagnetic, weak, strong, and Higgs) interactions with detectors. The predictions of the present theory include (1) the dark matter candidate of our previous papers, (2) many new fermions with masses not far above 1 TeV, and (3) the full range of superpartners with a modified phenomenology.
I Introduction
There are at least three compelling physical arguments for some form of supersymmetry Baer-Tata ; Nath ; Mohapatra ; Haber-Kane ; Kane-susy ; Martin ; Drees ; Baer-Barger-Tata ; Roszkowski-2018 ; Baer-Barger-2020 ; Tata-2020 ; Arkani-Hamed-Giudice ; Ellis-Olive ; 3 : The Higgs mass can be protected from radiative corrections that would lift it to absurdly large values; unification of nongravitational coupling constants can be achieved at high energy; and, assuming R-parity conservation, the lightest superpartner (LSP) should be a dark matter constituent.
An additional argument is aesthetic rather than physical: the mathematical beauty of conventional supersymmetry (SUSY), with a super-Poincaré algebra where
$\left\{Q_{\alpha},\overline{Q}_{\dot{\beta}}\right\}=2\left(\sigma^{\mu}\right)_{\alpha\dot{\beta}}P_{\mu}$ (1)
(in a standard notation), and the elegance of its extensions up to supergravity, string theory, and beyond.
However, Nature is not required to respect human aesthetic preferences, and we are now at a point where there is considerable skepticism about the the viability of any version of supersymmetry, following the exclusion at the LHC of expected superpartners in the most favored range of masses. Some skepticism was already being expressed 25 years ago, even by speakers at supersymmetry conferences, but the lack of evidence for SUSY is now often perceived to represent a crisis in fundamental physics.
In addition to this tension of SUSY with experiment, there are also theoretical impediments. For example, breaking SUSY is a difficult problem for which no convincing solution has been found.
In this paper we consider an alternative and highly unconventional form of supersymmetry – called susy here to avoid confusion – which is compatible with the physical motivations listed above, but which redirects aesthetic preferences: There is a superpartner for each standard model particle, the Higgs mass is protected as usual from a nonlogarithmic divergence at high energy, unification of nongravitational coupling constants can be achieved, and the LSP is a dark matter constituent. This last feature, however, is now placed in a multi-component scenario, with the nonsupersymmetric WIMP of our previous papers DM2021a ; DM2021b ; DM2022a ; DM2022b ; perspective assumed to be the dominant constituent because of its relatively low mass.
As will be seen below, the present formulation leads to reduced cross-sections for the production and decay of superpartners, and to modified experimental signatures, which may explain why none have yet been identified.
Furthermore, in the present formulation susy is automatically broken by the requirement of a stable vacuum (or, equivalently in this context, Lorentz invariance) as will again be seen below.
The breaking of conventional SUSY has been a central and unsolved problem for about half a century Witten . The difficulty ultimately results from (1), which implies that a mechanism for breaking conventional susy must increase the energy Nath ; Witten , in contrast to the normal symmetry breakings elsewhere in physics which lower the energy.
In the present formulation, the unsolved “Gordian knot” of supersymmetry breaking (and related problems) is severed by postulating a simple primitive susy with no counterpart to (1), in which each standard model Weyl fermionic field $\psi^{r}_{f}$ is matched with a primitive (unphysical) bosonic field $\psi^{r}_{b}$;
$\psi^{r}=\left(\begin{array}[]{c}\psi^{r}_{b}\\ \psi^{r}_{f}\end{array}\right)\quad\quad\mathrm{or}\quad\quad\psi=\left(\begin{array}[]{c}\psi_{b}\\ \psi_{f}\end{array}\right)\;.$ (6)
Here $\psi^{r}_{f}$ and $\psi^{r}_{b}$ are both 2-component spinors, with both belonging to the same gauge representation and having the same quantum numbers; $\psi_{f}$ and $\psi_{b}$ represent the full sets of these fields corresponding to all physical gauge representations. There is then a trivial invariance of the initial action
$S_{\mathrm{prim}}$ $=S_{f}+S_{b}$ (7)
$S_{f}=\int d^{4}x\,e\,\overline{\mathcal{L}}_{f}\quad$ ,¯$,\quad\overline{\mathcal{L}}_{f}=\psi_{f}^{\dagger}ie_{\alpha}^{\mu}\sigma^{\alpha}D_{\mu}\psi_{f}$ (8)
$S_{b}=\int d^{4}x\,e\,\overline{\mathcal{L}}_{b}\quad$ ,¯$,\quad\overline{\mathcal{L}}_{b}=\psi_{b}^{\dagger}ie_{\alpha}^{\mu}\sigma^{\alpha}D_{\mu}\psi_{b}$ (9)
(again in a standard notation) under the simplest of rotations of $\psi^{r}$, using Grassmann parameters $\epsilon$ and $\epsilon^{{\dagger}}$.
The present formulation thus trivially satisfies the Coleman-Mandula theorem Coleman-Mandula , which implies that the symmetries described by ordinary Lie algebras (with only bosonic elements) cannot mix fields of different spins. In conventional SUSY, this theorem is evaded via the extension to a graded algebra with both bosonic and fermioinc elements. Here it is satisfied because the primitive susy transformations involving the primitive bosonic fields $\psi_{b}$ manifestly do not mix different spin states.
The $\psi_{b}$ are grossly unphysical, because their excitations would have negative energy (and, of course, they violate the spin-statistics theorem required by Lorentz invariance).
Physically acceptable bosonic fields (with excitations $a^{{\dagger}}\ket{0}$ having positive energy) are obtained by transforming the primitive $\psi_{b}$ in exactly the same way as in our previous papers, which omitted susy but led to a dark matter candidate with well-defined mass and couplings DM2021a ; DM2021b ; DM2022a ; DM2022b ; perspective . Just as before we will find that the final result is a reinterpretation of the scalar boson sectors: The original degrees of freedom in the $\psi_{b}$ are transformed into three kinds of scalar boson fields: complex Higgs-like fields $\phi$, auxiliary fields $F$, and real fields $\varphi$ of a new kind.
The treatment below is thus an extension of the same phenomenology as in Ref. DM2021a , but there is an important difference: In the Higgs-related sectors, both Higgs and higgson mass eigenstates can be formed from the initial degrees of freedom, as described below, since there are two fields available with the same standard-model quantum numbers. But within the context of $SO(10)$ grand unification Nanopoulos-1984 ; Barger and the present theory, the superpartners of standard model fermions – sfermions as redefined here – must be excitations of the real fields $\varphi$, with only the second-order interactions of (63)-(66). This implies reduced cross-sections and a profound effect on the experimental signatures.
In the next section this point is described in more detail, via a straightforward generalization of the arguments in Ref. DM2021a and perspective . The experimental consequences are discussed in the last section.
II Sfermions redefined
The primitive fields and action of (6)-(9) are derived in perspective , and shown in its (5.14) and (5.15), but are here simply postulated as the starting point of the present phenomenology.
A complicated but straightforward set of transformations, given in Section VII and Appendix F of perspective , then results in the Lagrangian in (4) of DM2021a or (7.10) of  perspective for a gauge multiplet $\Phi$ consisting of partially transformed but still primitive and unphysical bosonic fields $\Phi_{r}$, before the addition of masses and interactions:
¯$\overline{\mathcal{L}}_{\Phi}=\Phi^{{\dagger}}\left(x\right)D^{\mu}D_{\mu}\Phi\left(x\right)+\Phi^{{\dagger}}\left(x\right)B_{k}\,\sigma^{k}\,\Phi\left(x\right)\hskip 8.5359pt$ (10)
where $D_{\mu}=\partial_{\mu}-iA_{\mu}$, $A_{\mu}=A_{\mu}^{i}t^{i}$, the $t^{i}$ are gauge-field generators, the $\sigma^{k}$ are Pauli matrices,
$F_{kk^{\prime}}=-\varepsilon_{kk^{\prime}k^{\prime\prime}}B_{k^{\prime\prime}}\quad\left(\mathrm{with}\;k,k^{\prime}=1,2,3\right),$ (11)
and $F_{\mu\nu}$ is the field strength tensor (with $\mu,\nu=0,1,2,3$).
The conventions of DM2021a (or perspective ) are retained here initially, with a single unified gauge group and the coupling constant absorbed into $A_{\mu}$. Summations are implied over coordinate and gauge-field indices like $\mu$, $k$, and $i$, but not labels of other fields, like $r$, $R$, $s$, $S$, $I$, and $\alpha$. It is straightforward to adapt the treatment of this section to the gauge groups of the standard model after symmetry breaking, and this is done in the last section. Auxiliary fields $F_{R}$ also automatically emerge from these transformations, with the Lagrangian in (9.2) of perspective :
¯$\overline{\mathcal{L}}_{F}=\sum_{R}F_{R}^{\dagger}\left(x\right)F_{R}\left(x\right)\;.$ (12)
The excitations of physical boson fields must have positive energy, a requirement closely related to Lorentz invariance. This can be achieved in two steps: (i) The anomalous second term of (10) must be eliminated, and (ii) the original 2-component spinor fields $\Phi$ must be reduced to 1-component scalar fields $\phi$ or $\varphi$, as defined below.
Conventional scalar boson fields $\phi$ can be constructed by combining $2$-component fields $\Phi_{r}$ and $\Phi_{r^{\prime}}$ with the same gauge quantum numbers but opposite spins (and equal amplitudes):
$\Phi_{R}=\left(\begin{array}[]{c}\Phi_{r}\\ \Phi_{r^{\prime}}\end{array}\right)$ (15)
so that
$\Phi_{R}^{{\dagger}}\left(x\right)\,\sigma^{k}\,\Phi_{R}\left(x\right)$ $=\left(\begin{array}[]{cc}\Phi_{r}^{{\dagger}}&\Phi_{r^{\prime}}^{{\dagger}}\end{array}\right)\left(\begin{array}[]{cc}\sigma^{k}&0\\ 0&\sigma^{k}\end{array}\right)\left(\begin{array}[]{c}\Phi_{r}\\ \Phi_{r^{\prime}}\end{array}\right)$ (21)
$=\Phi_{r}^{{\dagger}}\,\sigma^{k}\,\Phi_{r}+\Phi_{r^{\prime}}^{{\dagger}}\,\sigma^{k}\,\Phi_{r^{\prime}}$
$=0\;.$ (22)
We can then choose
$\Phi_{R}\left(x\right)=\phi_{R}\left(x\right)\,\xi_{R}\quad\mathrm{\ with}\quad\xi_{R}^{\,{\dagger}}\,\xi_{R}=1$ (23)
where $\xi_{R}$ has $4$ constant components and $\phi_{R}\left(x\right)$ is a $1$-component complex amplitude which is roughly analogous to the Higgs/amplitude modes predicted and observed in superconductors (and possibly other systems) Varma1 ; Varma2 ; Shimano .
With the generators in $D_{\mu}$ treated as operators, (22) and (23) imply that we can combine the original Lagrangians of (10) for $\Phi_{r}$ and $\Phi_{r^{\prime}}$ to obtain
¯$\overline{\mathcal{L}}_{R}=\phi_{R}^{\ast}\left(x\right)D^{\mu}D_{\mu}\,\phi_{R}\left(x\right)$ (24)
or for a gauge multiplet $\widetilde{\phi}_{\widetilde{R}}$
¯$\overline{\mathcal{L}}_{\widetilde{R}}=\widetilde{\phi}_{\widetilde{R}}^{{\dagger}}\left(x\right)D^{\mu}D_{\mu}\widetilde{\phi}_{\widetilde{R}}\left(x\right)\;.$ (25)
We thus recover the usual result for conventional scalar boson fields $\phi_{R}$ – i.e. Higgs and conventional sfermion fields – if these fields can be formed through the above procedure.
This is possible for the usual electroweak Higgs field within conventional theories. For example, within an $SO(10)$ grand-unified theory there are various doublets with the same relevant quantum numbers, together with their charge conjugates Pernow .
However, as will be seen below, this is not possible for the conventional sfermion fields.
Unconventional scalar boson fields $\overline{\varphi}$ (and later $\varphi$) can be constructed by combining a $2$-component field $\Phi_{s}$ and a field $\Phi_{s}^{c}$ which is charge-conjugate to $\Phi_{s}$, in the sense that it has opposite gauge quantum numbers, but is here taken to have the same spin as well as the same amplitude:
$\Phi_{S}=\left(\begin{array}[]{c}\Phi_{s}\\ \Phi_{s}^{c}\end{array}\right)\quad\mathrm{with}\quad\Phi_{s}^{c\,{\dagger}}\Phi_{s}^{c}=\Phi_{s}^{{\dagger}}\Phi_{s}\;.$ (28)
The subscripts $s$ and $S$ are used here to avoid confusion. In the present context (and most obviously for charged particles) $\Phi_{s}$ and $\Phi_{s}^{c}$ must exist independently in the initial Lagrangian. It will be seen below that this implies the existence of additional fermions awaiting discovery, with masses not extremely far above 1 TeV.
Since
$\Phi_{s}^{c\,{\dagger}}\,t^{i}\,\Phi_{s}^{c}=-\,\Phi_{s}^{{\dagger}}\,t^{i}\,\Phi_{s}$ (29)
where the generators $t^{i}$ are again treated as operators, and
$B_{k^{\prime\prime}}=-\varepsilon_{k^{\prime\prime}kk^{\prime}}F_{kk^{\prime}}\quad,\quad F_{kk^{\prime}}=F_{kk^{\prime}}^{j}t^{j}\;,$ (30)
we have
$\Phi_{S}^{{\dagger}}\left(x\right)B_{k}\,\Phi_{S}\left(x\right)$ $=\left(\begin{array}[]{cc}\Phi_{s}^{{\dagger}}&\Phi_{s}^{c\,{\dagger}}\end{array}\right)\left(\begin{array}[]{cc}B_{k}&0\\ 0&B_{k}\end{array}\right)\left(\begin{array}[]{c}\Phi_{s}\\ \Phi_{s}^{c}\end{array}\right)$ (36)
$=\Phi_{s}^{{\dagger}}B_{k}\,\Phi_{s}+\Phi_{s}^{c\,{\dagger}}B_{k}\Phi_{s}^{c}$
$=0\;.$ (37)
A similar cancellation and simplification occurs in
$\Phi_{S}^{{\dagger}}D^{\mu}D_{\mu}\Phi_{S}=$ $\Phi_{s}^{{\dagger}}\left(\partial^{\mu}-iA^{\mu i}t^{i}\right)\left(\partial_{\mu}-iA_{\mu}^{j}t^{j}\right)\Phi_{s}$
$\hskip 9.95863pt+\Phi_{s}^{c\,{\dagger}}\left(\partial^{\mu}-iA^{\mu i}t^{i}\right)\left(\partial_{\mu}-iA_{\mu}^{j}t^{j}\right)\Phi_{s}^{c}$ (38)
$=\Phi_{S}^{{\dagger}}\,\partial^{\mu}\partial_{\mu}\Phi_{S}-\Phi_{S}^{{\dagger}}\,A^{\mu}A_{\mu}\,\Phi_{S}$ (39)
with (29) used for each momentum eigenstate in $\Phi_{S}$, after $-i\partial_{\mu}\rightarrow p_{\mu}$.
For (28), (37), and (39) to be physically meaningful, it is necessary that both $\Phi_{s}$ and $\Phi_{s}^{c}$ are included in the multiplet (or multiplets) described by (10), so that the Lagrangian contains both of the terms
¯$\overline{\mathcal{L}}_{s}$ $=$ $\Phi_{s}^{{\dagger}}D^{\mu}D_{\mu}\Phi_{s}+\Phi_{s}^{{\dagger}}B_{k}\,\sigma^{k}\,\Phi_{s}$ (40)
¯$\overline{\mathcal{L}}_{s}^{c}$ $=$ $\Phi_{s}^{c\,{\dagger}}D^{\mu}D_{\mu}\Phi_{s}^{c}+\Phi_{s}^{c\,{\dagger}}B_{k}\,\sigma^{k}\,\Phi_{s}^{c}$ (41)
which will then reduce to
¯$\overline{\mathcal{L}}_{S}$ $=$ $\Phi_{S}^{{\dagger}}\,\partial^{\mu}\partial_{\mu}\Phi_{S}-\Phi_{S}^{{\dagger}}\,A^{\mu}A_{\mu}\,\Phi_{S}$ (42)
$=$ $\overline{\varphi}_{S}^{\ast}\,\partial^{\mu}\partial_{\mu}\overline{\varphi}_{S}-\overline{\varphi}_{S}^{\ast}\,A^{\mu}A_{\mu}\,\overline{\varphi}_{S}$ (43)
where
$\Phi_{S}\left(x\right)=\overline{\varphi}_{S}\left(x\right)\,\zeta_{S}\quad\mathrm{with}\quad\zeta_{R}^{\,{\dagger}}\,\zeta_{R}=1$ (44)
and $\overline{\varphi}_{S}$ is again a 1-component complex amplitude.
Notice that (43) and the steps leading to it are gauge invariant: If $A_{\mu}$ and (the mathematically fundamental but physically hidden) $\Phi_{s}$ are subjected to a gauge transformation, (43) still holds with the transformed $A_{\mu}$ and transformed $\overline{\varphi}_{S}$ of (44). The first-order terms have cancelled, and $\overline{\varphi}_{S}$ is transformed through the relations above rather than a simple phase factor, but the action of (43) is still invariant under a gauge transformation.
I.e., if the gauge fields are subjected to a gauge transformation, the fields $\overline{\varphi}$ are modified in a well-defined way which preserves the action, and this is what is meant by gauge invariance.
The final Lagrangians (24) and (43) for the physical scalar boson fields $\phi$ and $\overline{\varphi}$ are also manifestly Lorentz invariant.
As mentioned above, the procedure leading to the fields $\overline{\varphi}_{S}$ is meaningful only if the multiplet or multiplets containing a field $\Phi_{s}$ also contain the charge conjugate $\Phi_{s}^{c}$. This is satisfied in standard theories for the Higgs-like fields discussed in our previous papers – for example, with two doublets as in standard SUSY or with the $10\rightarrow 5+\overline{5}$ etc. of $SO(10)$. It is satisfied in the present theory for the sfermion-like fields when we include the full $16+\overline{16}$ fundamental spinor representation Ozer , for both fermion and primitive sfermion fields, rather than just the conventional $16$.
Fermions belonging to the $\overline{16}$ representation are then necessarily predicted, with masses not far from 1 TeV. (In the present theory all fields initially have the same chirality, so the $16$ and $\overline{16}$ are not equivalent. This is consistent with the standard treatment of $SO(10)$, in which all the physical fermion fields are initially left-handed.)
The primitive sfermion fields of the $16$ and $\overline{16}$ respectively provide the $\Phi_{s}$ and $\Phi_{s}^{c}$ required above to form physical but unconventional sfermion fields $\overline{\varphi}_{S}$ of a new kind, with only the second-order interactions of (43).
Conventional sfermion fields, with the conventional Lagrangian (24) or (25), cannot be formed because there is no second field $\Phi_{r^{\prime}}$ with equal quantum numbers (including flavor) to be matched with a given $\Phi_{r}$.
It is also important to include the couplings to Higgs fields. We recover essentially the same relationships as in conventional SUSY Martin ; Djouadi when each primitive $\Phi_{s}$ contributes a term with the usual form
¯$\overline{\mathcal{L}}^{H}_{s}=-y_{s}^{2}\,\Phi_{s}^{{\dagger}}H^{{\dagger}}H\Phi_{s}$ (45)
where $y_{s}$ is the Yukawa coupling constant for a Higgs field $H$ and the Weyl fermion field having the same quantum numbers as $\Phi_{s}$. (Complications associated with mass matrices are beyond the scope of the present paper.) The conjugate field $\Phi_{s}^{c}$ must similarly contribute
¯$\overline{\mathcal{L}}^{H}_{s\,c}=-y_{s\,c}^{2}\Phi_{s}^{c\,{\dagger}}H^{{\dagger}}H\Phi_{s}^{c}$ (46)
where $y_{s\,c}$ is the Yukawa coupling for the conjugate fermion field (in the as yet unobserved $\overline{16}$). The total interaction Lagrangian for each physical (complex) sfermion is therefore
¯$\overline{\mathcal{L}}^{H}_{S}$ $=-y_{S}^{2}\,\Phi_{S}^{{\dagger}}H^{{\dagger}}H\Phi_{S}=-y_{S}^{2}\,\overline{\varphi}_{S}^{\ast}\,H^{{\dagger}}H\overline{\varphi}_{S}$ (47)
$y_{S}^{2}$ $=\frac{1}{2}\left(y_{s}^{2}+y_{s\,c}^{2}\right)$ (48)
(where the constraint of (28) has been used).
For a given Weyl fermion, there are two values of $s$, corresponding to spin up or down in the 2-component $\Phi_{s}$. As a result there are two complex sfermion fields $\overline{\varphi}_{S}$, each with a Higgs coupling constant of $\left(y_{s}^{2}+y_{s\,c}^{2}\right)/2$, to partner the two Weyl fermion fields with Yukawa couplings of $y_{s}$ and $y_{s\,c}$. A Dirac field, consisting of 2 Weyl fields, is effectively partnered with 2 complex sfermion fields, just as in conventional SUSY Martin ; Djouadi .
It may be appropriate to mention that the need to join two fields in forming a mass eigenstate, in order to satisfy symmetry requirements, is not unprecedented: The real and imaginary parts of an ordinary charged scalar field must be joined to achieve gauge invariance, and left-and right-handed Weyl fields must be joined to form a Dirac field $\psi_{D}$ or Majorana field $\psi_{M}$, in order to achieve Lorentz invariance.
III Experimental consequences
The treatment of the preceding section can be trivially modified to obtain the $SU(3)\times SU(2)\times U(1)$ version of the last term in (39), with coupling constants displayed:
¯$\overline{\mathcal{L}}^{int}_{S}$ $=-\sum_{n}g_{n}^{2}\,\Phi_{S}^{{\dagger}}\,A_{n}^{\mu i}\,t_{n}^{i}\,A_{n\mu}^{j}t_{n}^{j}\,\Phi_{S}\quad,\quad n=1,2,3$ (49)
$=-\frac{1}{2}\sum_{n}g_{n}^{2}\,A_{n}^{\mu i}\,A_{n\mu}^{j}\,\Phi_{S}^{{\dagger}}\left(t_{n}^{i}\,t_{n}^{j}+t_{n}^{j}\,t_{n}^{i}\right)\Phi_{S}\;.$ (50)
For the $SU(2)$ generators $\tau^{i}\rightarrow\sigma^{i}/2$ in a matrix representation, we have
$\sigma^{i}\sigma^{j}+\sigma^{j}\sigma^{i}=2\,\delta^{ij}$ (51)
and for the $SU(3)$ generators $T^{i}\rightarrow\lambda^{i}/2$
$\lambda^{i}\lambda^{j}+\lambda^{j}\lambda^{i}=\frac{4}{3}\delta^{ij}+2\,d^{ijk}\lambda^{k}$ (52)
or
$T^{i}T^{j}+T^{j}T^{i}=\frac{1}{3}\delta^{ij}+d^{ijk}T^{k}\quad\mathrm{with}\quad\Phi_{S}^{{\dagger}}T^{k}\Phi_{S}=0$ (53)
where the $d^{ijk}$ are structure constants. Then (50) reduces to
¯$\overline{\mathcal{L}}^{int}_{S}$ $=$ $-\sum_{n}\overline{g}_{n}^{2}\,\Phi_{S}^{{\dagger}}A_{n}^{\mu i}\,A_{n\mu}^{i}\Phi_{S}\quad,\quad n=1,2,3$ (54)
$=$ $-\sum_{n}\overline{g}_{n}^{2}\,\overline{\varphi}_{S}^{\ast}A_{n}^{\mu i}\,A_{n\mu}^{i}\overline{\varphi}_{S}$ (55)
with $\overline{g}_{3}^{2}=g_{3}^{2}/6$, $\overline{g}_{2}^{2}=g_{2}^{2}/4$, $\overline{g}_{1}^{2}=g_{1}^{2}$. The real and imaginary parts separate in (55), which can be written as the sum of the two terms
¯$\overline{\mathcal{L}}^{int}_{SI}=-\sum_{n}\overline{g}_{n}^{2}\,\varphi_{SI}^{{\dagger}}A_{n}^{\mu i}\,A_{n\mu}^{i}\varphi_{SI}\quad,\quad I=1,2$ (56)
where
$\overline{\varphi}_{S}=\varphi_{S1}+i\varphi_{S2}\;.$ (57)
Below we simplify the notation, letting
$\alpha\leftrightarrow S,I\;.$ (58)
When the original $SU(2)\times U(1)$ fields are rotated into those of the electroweak theory after symmetry breaking  Peskin
$W^{\pm}_{\mu}$ $=\frac{1}{\sqrt{2}}\left(A^{1}_{2\mu}\mp iA^{2}_{2\mu}\right)$ (59)
$Z_{\mu}$ $=\frac{1}{\sqrt{g_{1}^{2}+g_{2}^{2}}}\left(g_{2}A^{3}_{2\mu}-g_{1}A_{1\mu}\right)$ (60)
$\bar{A}_{\mu}$ $=\frac{1}{\sqrt{g_{1}^{2}+g_{2}^{2}}}\left(g_{1}A^{3}_{2\mu}+g_{2}A_{1\mu}\right)$ (61)
(56) gives for each real sfermion field
¯$\overline{\mathcal{L}}^{int}_{\alpha}=\overline{\mathcal{L}}^{QCD}_{\alpha}+\overline{\mathcal{L}}^{EM}_{\alpha}+\overline{\mathcal{L}}^{W}_{\alpha}+\overline{\mathcal{L}}^{Z}_{\alpha}$ (62)
with
¯$\overline{\mathcal{L}}^{QCD}_{\alpha}$ $=-\frac{g_{s}^{2}}{6}\,\varphi_{\alpha}^{{\dagger}}\,{\cal A}^{\mu i}{\cal A}_{\mu}^{i}\,\varphi_{\alpha}$ (63)
¯$\overline{\mathcal{L}}^{EM}_{\alpha}$ $=-\left(Qe\right)^{2}\,\varphi_{\alpha}^{{\dagger}}\,\bar{A}^{\mu}\bar{A}_{\mu}\,\varphi_{\alpha}$ (64)
¯$\overline{\mathcal{L}}^{W}_{\alpha}$ $=-\frac{g^{2}}{2}\,\varphi_{\alpha}^{{\dagger}}\,W^{+\mu}W^{-}_{\mu}\,\varphi_{\alpha}$ (65)
¯$\overline{\mathcal{L}}^{Z}_{\alpha}$ $=-\frac{g_{Z}^{2}}{4}\,\varphi_{\alpha}^{{\dagger}}\,Z^{\mu}Z_{\mu}\,\varphi_{\alpha}$ (66)
where $g_{s}=g_{3}$, $g=g_{2}$, $g_{Z}=g/\cos\theta_{W}$, $\cos\theta_{W}=g_{2}/\sqrt{g_{1}^{2}+g_{2}^{2}}$, $Qe$ is the electric charge, $\bar{A}_{\mu}$ is the electromagnetic vector potential, and ${\cal A}^{\mu}={\cal A}^{\mu i}T^{i}$ is the QCD gauge potential with gluon fields ${\cal A}_{\mu}^{i}$.
Corresponding to each Weyl fermion field $\psi$, there are initially two 2-component complex fields $\Phi_{s}$ (with opposite spins) having the same quantum numbers (including flavor) as $\psi$. These primitive (and unphysical) bosonic fields, with 4 total degrees of freedom, yield 4 real physical fields $\varphi_{\alpha}$ through the procedure described above. The $\varphi_{\alpha}$ (as mass eigenstates) are the final redefined sfermion fields of the present theory. They will have only the second-order gauge interactions of (64)-(66) and the second-order Higgs interactions of (48), which can now be rewritten
¯$\overline{\mathcal{L}}^{H}_{\alpha}=-y_{\alpha}^{2}\,\varphi_{\alpha}\,H^{{\dagger}}H\,\varphi_{\alpha}\;.$ (67)
As noted above, and indicated in Fig. 1, the interactions of (67) are exactly what is needed to give the usual susy cancellation of the quadratically-divergent radiative correction of quarks to the squared mass of the observed Higgs boson, reducing it to a manageable logarithmic divergence: For two conjugate Weyl fermions in the $16$ plus $\overline{16}$, there are four real sfermion fields $\varphi_{\alpha}$. They will have the same combined effect as two complex conventional sfermion fields, each with $y_{\alpha}^{2}=\frac{1}{2}\left(y_{s}^{2}+y_{s\,c}^{2}\right)$.
This cancellation persists to all orders, since the processes shown for Higgs bosons in Fig. 1 hold as well for gauge bosons in all virtual processes, with each vertex involving a boson and pair of fermions or else two pairs of bosons, as in the figure. There are still the usual logarithmic divergences, but if the most relevant superpartners are not far above 1 TeV the corrections to the Higgs mass are physically reasonable.
Many of the standard processes for squark and gluino production do not exist in the present scenario, because they involve first-order interactions, as illustrated in Fig. 2. See also Fig. 14 etc. of Haber-Kane , or the corresponding figures in Baer-Tata , Drees , and many other books and papers.
It should be emphasized, however, that there are other processes in conventional SUSY (and shown in publications like these) which persist in the present description, with the same amplitudes as originally calculated. For example, production of squarks by direct gluon fusion and higher-order vector-boson fusion is still allowed, with the first of these depicted in Fig. 1. Calculations in our group by Tallman et al. Tallman show that the total production cross-section at the 13 TeV LHC is relatively small for a squark mass of 400 GeV– about 1 pb – and decreases with higher masses (although increasing by orders of magnitude at substantially lower masses).
The phenomenology of more general superpartner production and decay will also be modified, again because processes requiring first-order sfermion interactions (as in e.g. Fig. 33 (b) and (c) of Haber-Kane ) will not exist, and this will reduce the total cross-sections for production and decay. In fact, the most readily detectable conventional processes for decay of gluinos and other superpartners are largely forbidden. An example is shown in Fig. 3.
Figure 1: Left panel: Representative diagrams for contributions of a fermion and sfermion – in this case top quark and top squark – to quantum corrections of Higgs mass-squared. According to (67), the unconventional sfermions defined here will still provide the standard susy cancellation of quadratric divergences, provided that all fermions and (redefined) sfermions have masses $\lesssim$ a few TeV. As noted in the text, this cancellation holds for all processes in higher-order diagrams, with fermions and sfermions coupled to both Higgs bosons and gauge bosons. There will still be logarithmic corrections from fermions and sfermions, just as in conventional SUSY, but this is consistent with experiment for the particles with masses not far above 1 TeV implied by the present theory. Right panel: Most squark production processes are not allowed in the present description, because they would require first-order interactions, but production by direct gluon fusion and higher-order vector-boson fusion is still possible.
Figure 2: With the present interpretation of sfermions, squarks $\widetilde{q}$ cannot be produced in processes like these, involving quarks $q$ and gluons $g$ in colliding protons, since each vertex (i.e. interaction) in a physical process must involve two squark fields and two other bosonic fields.
Figure 3: Left panel: All sfermion decay processes, like the one shown here, do not exist (again because each vertex in a physical process must involve two sfermion fields and two other bosonic fields). The standard schemes for detecting sfermions rely on their being produced in collisions through processes that mostly do not exist in this new interpretation, and then decaying through processes that are entirely disallowed in this interpretation. This implies that new detection schemes are required, and that squarks with masses $\sim$ 1 TeV may exist even though they have not previously been identified. Right panel: Example of a process after gluino production that will not be seen in the present scenario, because it requires first-order squark vertices.
The squarks defined here can radiate pairs of gauge bosons (gluons, $W$ and $Z$ bosons, and photons) via (63)-(66), and in principle even Higgs bosons through (67). They cannot decay, however, because of the form of these interactions in (63)-(67): A single incident squark must always emerge after any process. The same is true of the other sfermions defined here.
The search for supersymmetry is very well motivated  Baer-Tata ; Mohapatra ; Nath ; Haber-Kane ; Kane-susy ; Drees ; Baer-Barger-Tata ; Roszkowski-2018 ; Baer-Barger-2020 ; Tata-2020 ; Arkani-Hamed-Giudice ; Ellis-Olive ; 3 , and it has been a major mystery that not a single superpartner has yet been discovered. Within the present description, this may be because the cross-sections for both production and decay are reduced, and the phenomenology is very different from what has been expected (and incorporated in the analysis and simulation of events).
One principal prediction is that sfermions may still be produced through direct gluon-gluon fusion, or higher-order vector-boson fusion, but cannot be observed through either decay products or the usual signatures resulting from first-order interactions. Instead, they can cause transitions in detector components via second-order (electromagnetic, weak, strong, and Higgs) interactions. I.e., they can still be directly detected via straight tracks left by these interactions, even though they have no net charge of any kind, and this is another aspect of the modified phenomenology. In addition, they can be indirectly detected via radiation of gauge bosons.
Gauginos and higgsinos can be observed somewhat in the usual way, but will have reduced cross-sections for production and decay because many processes involving squarks are disallowed. Calculations of cross-sections etc. for the remaining allowed processes are beyond the scope of the present paper but are planned Tallman .
Since sfermions in the present picture can annihilate but not decay, they are predicted to be very minor components of the dark matter: They are stable, electrically neutral, and colorless (with zero expectation value for every charge operator), as required for dark matter. But in the early universe, with larger masses and typically stronger interactions, they will have a lower thermal abundance, and in most cases will also annihilate more rapidly, so that their relic abundances will be much lower than that of a 70 GeV WIMP DM2021a ; DM2021b ; DM2022a ; DM2022b as they are thinned out in an expanding universe.
The present theory leads to several predictions that are testable within the foreseeable future: (1) the dark matter candidate of Refs. DM2021a ; DM2021b ; DM2022a ; DM2022b , with well-defined mass and detection cross-sections; (2) a large number of new fermions with masses not far from 1 TeV, which result from the full fundamental $16+\overline{16}$ representation of $SO(10)$ being physically realized, rather than just the conventional $16$; and (3) a new phenomenology for squarks, sleptons, gauginos, and higgsinos (and their linear combinations), providing a potential explanation of the fact that superpartners have not yet been identified.
Acknowledgement, competing interests, and data availability
The Feynman diagrams of Figs. 1-3 were prepared by Bailey Tallman. The author has no competing interests. No new data was used in this work.